Implement the following function using two op-amps and the required number of resistors

Vo = V1 - 2V2 + 3V3 - 4V4

1
Expert's answer
2021-05-31T06:18:33-0400


Applying the KCL at node 1

"\frac{0-V_1}{1K\Omega}-\frac{0-2V_2}{1K\Omega}+\frac{0-V_{01}}{1K\Omega}=0"

"V_{01}=-V_1-2V_2"

Applying the KCL at node 2

"\frac{0-V_1}{1K\Omega}+\frac{0-3V_3}{1K\Omega}+\frac{0-4V_{4}}{1K\Omega}+\frac{0-V_{0}}{1K\Omega}=0"

"-V_{01}-3V_3-4V_4=V_0"

"V_O=-[-V_1-2V_2]-3V_3-4V_4"

"V_O=V_1+2V_2-3V_3-4V_4"